OpenStudy (pphalke):

find an equation of the line that satisfies the given conditions a) standard form b) slope intercept form x intercept is (9,0) and slope: 8

OpenStudy (pphalke):

@math&ing001 @SnuggieLad @ONYXxELITE

OpenStudy (xguardians):

Do you know the equation for the standard form of a line?

OpenStudy (mathmale):

Please find and copy down here the "standard form" and the "slope-intercept" form of the equation of a straight line. You need to know what you're looking for in this problem.

OpenStudy (pphalke):

standard form is Ax+By=C and slope intercept is y=mx+b

OpenStudy (pphalke):

i found that the slop intercept for the above equation is y=9x+8 am i right?

OpenStudy (xguardians):

No that equation is incorrect.

OpenStudy (pphalke):

how is it supposed to be?

OpenStudy (xguardians):

First you must find the b value by plugging in the coordinate it gives you. y = mx + b We know the slope is 8 y = 8x + b Plug in the coordinate it gave you. 0 = 8(9) + b 0 = 72 + b b = -72 y = 8x - 72

OpenStudy (pphalke):

@xGuardians did u read wat the question wants?

OpenStudy (xguardians):

Yes. I just gave you the answer to the point-slope equation (b).

OpenStudy (mathmale):

Your goal: Find the line such that "x intercept is (9,0) and slope: 8" This means that when x=8, y=0. xGuardians' work is correct; what makes you think he didn't read the instructions?

OpenStudy (pphalke):

and in standard form?

OpenStudy (mathmale):

Take the equation arrived at by xGuardians and convert it into standard form. You correctly identified the appearance of "standard form."

OpenStudy (mathmale):

In other words, convert y = 8x - 72 into standard form Ax+By=C.

OpenStudy (pphalke):

wouldnt it be 8x-72+9

OpenStudy (mathmale):

No, because this expression of yours has no = sign. As you yourself have stated, standard form looks like Ax+By=C. Where's your = sign?

OpenStudy (mathmale):

and what happened to the variable y?

OpenStudy (pphalke):

i dont know

OpenStudy (mathmale):

Back to the drawing board: y = 8x - 72 Note that this includes a = sign. We want both x and y to be on the same side of your equation. Please subtract 8x from both sides of y = 8x - 72.

OpenStudy (pphalke):

i get -80?

OpenStudy (mathmale):

I don't see how you'd get -80. Please note that you're working with an equation y = 8x - 72 x and y are NOT going to disappear. Basically you're just rearranging these 3 terms. If y = 8x - 72 please subtract 8x from both sides. Note that -72-8x is not -80; one term (-72) is just a number, whereas the other term (-8x) contains the variable x.

OpenStudy (pphalke):

how do i do that?

OpenStudy (mathmale):

Subtract 8x from both sides of your equation: y = 8x - 72 -8x = -8x ------------------ Finish this, please.

OpenStudy (mathmale):

Please do not move on to a new problem without finishing the current one. You are not finished with this problem. It is essential that you understand what "subtract 8x from both sides" means and how to do it.

OpenStudy (pphalke):

oh ok y=x-72?

OpenStudy (mathmale):

Subtract 8x from both sides of y = 8x - 72. No, it is not y=x-72.

OpenStudy (mathmale):

y = 8x - 72 -8x = -8x --------------------- Finish this, please.

OpenStudy (pphalke):

-8x+y=72?

OpenStudy (pphalke):

@mathmale

OpenStudy (mathmale):

Correct, except for the sign of 72. Why do you believe the sign of 72 is positive when it started out as negative?

OpenStudy (pphalke):

because as we have subtracted -8x the sign immediately changed